# Enabling Systematic Generalization in Abstract Spatial Reasoning through Meta-Learning for Compositionality

### v1

@@ Abstract
Systematic generalization refers to the capacity to understand and generate novel combinations from known components. Despite recent progress by large language models (LLMs) across various domains, these models often fail to extend their knowledge to novel compositional scenarios, revealing notable limitations in systematic generalization. There has been an ongoing debate about whether neural networks possess the capacity for systematic generalization, with recent studies suggesting that meta-learning approaches designed for compositionality can significantly enhance this ability. However, these insights have largely been confined to linguistic problems, leaving their applicability to other tasks an open question. In this study, we extend the approach of meta-learning for compositionality to the domain of abstract spatial reasoning. To this end, we introduce SYGAR—a dataset designed to evaluate the capacity of models to systematically generalize from known geometric transformations (e.g., translation, rotation) of two-dimensional objects to novel combinations of these transformations (e.g., translation+rotation). Our results show that a transformer-based encoder-decoder model, trained via meta-learning for compositionality, can systematically generalize to previously unseen transformation compositions, significantly outperforming state-of-the-art LLMs, including o3-mini, GPT-4o, and Gemini 2.0 Flash, which fail to exhibit similar systematic behavior. Our findings highlight the effectiveness of meta-learning in promoting systematicity beyond linguistic tasks, suggesting a promising direction toward more robust and generalizable models.
@@ 1 Introduction
Despite its central role in human cognition, systematic generalization remains a significant challenge in artificial intelligence (Lake & Baroni, 2018; Loula et al., 2018; Hupkes et al., 2020). While large language models have recently demonstrated notable progress across various domains (OpenAI, 2024; Guo et al., 2025), they often fail to combine acquired knowledge in novel scenarios, demonstrating notable difficulties with systematic generalization (Dziri et al., 2023; Ismayilzada et al., 2025; Gendron et al., 2024). The question of whether neural networks can achieve systematicity has been the subject of extensive debate (Fodor & Pylyshyn, 1988; Brakel & Frank, 2009; Calvo & Symons, 2014, inter alia). Recent research by Lake & Baroni (2023) demonstrates that a transformer-based encoder-decoder model, trained using meta-learning for compositionality (MLC), can achieve human-like systematic generalization in processing instructions expressed in a pseudolanguage. By training the model to combine basic units of pseudolanguage into novel sequences over a stream of dynamically changing grammars, Lake & Baroni (2023) show that this model can effectively generalize to previously unseen compositions of language (see Section 2 for further details). While this approach presents a promising direction for addressing systematicity in neural networks, its applicability beyond linguistic contexts remains an open question.
In this study, we extend the MLC framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. Drawing inspiration from the abstraction and reasoning corpus (Chollet, 2019), we introduce a new dataset that assesses systematic generalization with respect to basic geometric transformations of two-dimensional objects within grid-based environments (see Figure 1). Using MLC, we train an encoder-decoder model on the task and demonstrate that it generalizes to unseen compositions of geometric transformations, significantly outperforming state-of-the-art LLMs, including GPT-4o (Achiam et al., 2023), o3-mini (OpenAI, 2025), and Gemini 2.0 Flash (DeepMind, 2024), which fail to exhibit comparable systematic behavior. To the best of our knowledge, this is the first application of MLC in abstract spatial reasoning. In summary, our contributions are as follows:
@@
We introduce SYGAR—a new dataset designed to assess systematic generalization in abstract spatial reasoning. This dataset includes basic geometric transformations (e.g., translation, rotation) and their compositions (e.g., translation + rotation) applied to two-dimensional objects in a grid-based environment.
@@
We show that MLC enables models to generalize to unseen compositions of geometric transformations, highlighting its potential beyond linguistic tasks.
@@
We find that models trained via MLC significantly outperform state-of-the-art LLMs such as GPT-4o, o3-mini, and Gemini 2.0 Flash on this task.
@@ 2 Background: meta-learning for compositionality
While Lake & Baroni (2023) also evaluate MLC on COGS (Kim & Linzen, 2020) and SCAN (Lake & Baroni, 2018), which test systematic lexical generalization to novel word combinations, their experiments are confined to the linguistic domain. In the following section, we show how MLC can be extended to support systematic generalization in abstract spatial reasoning, demonstrating its potential beyond linguistic tasks.
@@ 3 Method
3.1 SYGAR: a dataset for SYstematic Generalization in Abstract spatial Reasoning
To test systematicity in abstract spatial reasoning, we leverage the closure property of combined geometric transformations, where the composition of two valid transformations—such as translation, rotation, and reflection—yields another valid geometric transformation (Brannan et al., 2011). Drawing inspiration from the abstraction and reasoning corpus (Chollet, 2019), we design a task in which abstract objects, defined in a two-dimensional grid environment, are subjected to basic geometric transformations and their compositions (see Figure 1 for examples). We use fixed-size grids, each of which can be represented as a two-dimensional array of integers, where different values correspond to distinct colors. We use integers from 0 to 9, with 0 denoting a black background and the remaining integers mapping to unique colors (see Appendix A.1 for more details). Objects are defined based on color connectivity; that is, each object comprises a group of connected cells sharing the same color. Connectivity is determined by the Moore neighborhood (Bays, 2010), meaning that cells are considered connected if they are directly or diagonally adjacent. Each grid contains either one or two objects. A transformation is represented as a pair of grids, with the input grid displaying the objects before, and the output grid showing them after the geometric transformation. Each transformation affects only one of the objects in the grid. For example, in Figure 1(a), a single L-shaped yellow object is translated one step downward. In Figure 1c, a square blue object in the bottom-right expands toward the neighboring top row. Objects never occlude one another nor extend beyond the boundaries of the grids.
@@ 3.1 SYGAR: a dataset for SYstematic Generalization in Abstract spatial Reasoning
To test systematicity, we present examples of primitive transformations and their compositions, and evaluate models on previously unseen combinations of indicators. For instance, in Figure 3, models are asked to infer the correct transformation for a previously unseen level-2 composition of indicators, given study examples of primitive transformations and their level-1 compositions. Conceptually, our setup is similar to the few-shot compositional task introduced by Lake & Baroni (2023) (see Section 2), but it replaces the lexical interpretation grammar with a visual interpretation grammar. Specifically, models need to infer which indicator maps to which transformation, and how to compose them to deduce the correct final transformation. For a detailed description of how we algorithmically generate dataset samples, please refer to Appendix A.3.
@@ 3.2 Meta-Learning for compositionality in abstract spatial reasoning
To systematically generalize from known geometric transformations to previously unseen transformation compositions, we extend the meta-learning for compositionality (Lake & Baroni, 2023) framework described in Section 2. As in the original MLC approach, we train a transformer-based encoder-decoder model on a dataset of dynamically changing interpretation grammars. However, instead of mapping pseudolinguistic instructions to sequences of abstract symbols, we consider a visual interpretation grammar that associates visual indicators (object shape, color, or proximity to an indicator object) with specific geometric transformations, as described in Section 3.1. An episode is defined as a set of study examples that demonstrate the underlying grammar, along with query inputs for which the correct outputs must be inferred (see Figure 3 for an illustration). By training over a series of episodes with changing visual interpretation grammars, the model needs to abstract and recombine information from the study examples in order to predict the correct output for previously unseen query grids.
@@ 4.1 Task setup
We consider two different task setups in this work. The first is a standard few-shot learning scenario, in which models are required to generate an output grid for a given query input that undergoes a level-2 transformation composition. This prediction is based on three examples that demonstrate the same level-2 transformation. A visual representation of this setup is provided in Figure 4 in the Appendix. This task evaluates the model’s ability to infer geometric transformations from a limited set of illustrative examples.
@@
The second setup focuses on systematicity and differs from the first in the type of few-shot examples presented. As mentioned in Section 3.1, the idea is to test whether models can infer novel compositions from known geometric transformations. For this, we replace the level-2 few-shot examples with a set of primitive transformations plus level-1 transformation compositions, and ask the model to predict the previously unseen level-2 transformation composition, as illustrated in Figure 3. Specifically, we present six primitive transformations—two examples for each indicator (shape-based, color-based, neighbor-based)—and six level-1 transformation compositions, two examples for each first-level indicator composition (shape+color, shape+neighbor, color+neighbor).
@@
We generate 100,000 episodes, each comprising three few-shot examples for the standard few-shot learning task, 12 systematic study examples for the systematicity setup, and ten query input-output grid pairs demonstrating the final level-2 transformation composition. Each episode is characterized by a unique visual interpretation grammar. For instance, in one episode, yellow objects are translated downward by a single cell, while in another, yellow objects are reflected horizontally. This challenges models to generalize to novel scenarios by inferring the correct transformations based on the provided examples. To train our encoder-decoder model via MLC, we split the data into 82,908 training, 8,546 validation and 8,546 test episodes. Importantly, the data splits are constructed such that the geometric transformations involved in the final query level-2 compositions differ between the training and evaluation sets. For instance, while the model is trained on all basic transformations and a series of transformation compositions (e.g., translation+rotation+reflection), it is tested on compositions not seen during training (e.g., translation+rotation+extension). For comprehensive statistics of the dataset splits, please refer to Appendix A.4.
4.2 Language models
@@
In addition to the model trained via MLC, we evaluate three state-of-the-art LLMs on the test set of our proposed dataset: GPT-4o (Achiam et al., 2023), o3-mini (OpenAI, 2025), and Gemini 2.0 Flash (DeepMind, 2024). To textually prompt the models for a given episode, we represent grids as two-dimensional arrays, consistent with prior work (Moskvichev et al., 2023). We also test a multimodal setup in which both an image of the study examples and the input query are provided alongside the text prompt (text+image). Due to financial constraints, each model is evaluated on a single test query for each episode across the 8,546 episodes in the test set. All textual and visual prompts, specific model versions, and decoding parameters are detailed in Appendix C.2.
@@ 4.3 Evaluation metrics
To assess the quality of the generated output grids, we employ three different evaluation metrics: i) accuracy, ii) color accuracy, and iii) shape accuracy. For accuracy, we use exact match, i.e., an output is considered accurate if all cells in the predicted output grid correspond to those in the target grid. Color accuracy measures whether the predicted objects in the output grid match the colors of the objects in the target grid, regardless of their shape or location. Shape accuracy, on the other hand, evaluates whether the predicted objects share the same shape as those in the target grid, irrespective of color or location. For formal definitions of these metrics, please refer to Appendix C.1.
| Model | Accuracy [%] | Color Acc. [%] | Shape Acc. [%] | |
| 3-Shot | GPT-4o (text-only) | |||
| GPT-4o (text+image) | ||||
| Gemini 2.0 Flash (text-only) | ||||
| Gemini 2.0 Flash (text+image) | ||||
| o3-mini | ||||
| MLC | 99.92 | 100.00 | 99.92 | |
| Systematicity | GPT-4o (text-only) | |||
| GPT-4o (text+image) | ||||
| Gemini 2.0 Flash (text-only) | 99.68 | |||
| Gemini 2.0 Flash (text+image) | ||||
| o3-mini | ||||
| MLC | 78.26 | 80.49 |
@@ 5 Results
In Table 1, we report the performance of our model trained via MLC, alongside the LLMs we evaluate on the two task setups, as described in Section 4.1.
@@
We begin by examining model performance on the three-shot learning task. In this setup, models are provided with three input-output examples that illustrate the final transformation (see Figure 4 in the Appendix for a visual illustration). Despite this guidance, LLMs such as GPT-4o and Gemini 2.0 Flash exhibit notable difficulties with this task. GPT-4o achieves a maximum accuracy of 22.28% (text-only), while Gemini 2.0 Flash performs marginally better with an accuracy of up to 30.08% (text-only). Although the long-chain-of-thought model o3-mini achieves a higher performance of up to 64.04%—performing best among all LLMs—its accuracy remains modest given the simplicity of the transformations involved. In contrast, our encoder-decoder model trained via MLC achieves an accuracy of up to 99.92%, substantially surpassing all other models. When comparing the models’ overall accuracy with their shape and color accuracy, we find that all models perform nearly perfectly in predicting object color. Notably, for GPT-4o and Gemini 2.0 Flash, shape accuracy is significantly higher than exact match accuracy. This discrepancy suggests that while these models are often able to predict the correct shape of an object, they frequently fail to accurately predict its final position. Interestingly, both GPT-4o and Gemini 2.0 Flash show a decline in accuracy when an additional visual input is included alongside the textual prompt. We hypothesize that this performance degradation stems from the added complexity introduced by the image, either due to modality alignment challenges (Masry et al., 2025) or limitations in leveraging the visual content for reasoning.
@@
In the systematicity task, models are asked to infer the correct final transformation composition from a set of study examples that represent more basic, decomposed transformations (see Figure 3 for an example). As shown in Table 1, all LLMs perform poorly on this task. For instance, GPT-4o achieves a maximum accuracy of 0.99% (text-only), while Gemini 2.0 Flash reaches up to 2.66% (text-only). Interestingly, o3-mini, the model that performed best among all LLMs on the standard few-shot learning task, performs worst in this setting, with an accuracy of only 0.53%. In contrast, our encoder-decoder model trained via MLC achieves an accuracy of up to 78.26%, significantly outperforming all other LLMs despite having orders of magnitude fewer parameters. Importantly, as described in Section 4.1, this model has never seen the specific level-2 compositions of geometric transformations in the test queries during training but was instead optimized on a distinct set of transformation compositions (see Table 5 in the Appendix). Consistent with our findings from the 3-shot learning task, models generally succeed in predicting the correct object colors. However, shape accuracy declines considerably in this setting, indicating that models struggle to predict the correct shape of the output object after the transformation. A qualitative example of the models’ predictions can be found in Figure 3, with additional examples provided in Appendix D. The strong performance of our model trained via MLC highlights the effectiveness of this training strategy in promoting systematic generalization to novel transformation compositions. The model not only learns to infer a visual interpretation grammar from a limited number of study examples but also generalizes to novel transformation compositions that it has never encountered during training.

### v2

# Compositional–ARC: Assessing Systematic Generalization in Abstract Spatial Reasoning

@@ Abstract
Systematic generalization refers to the capacity to understand and generate novel combinations from known components. Despite recent progress by large language models (LLMs) across various domains, these models often fail to extend their knowledge to novel compositional scenarios, revealing notable limitations in systematic generalization. There has been an ongoing debate about whether neural networks possess the capacity for systematic generalization, with recent studies suggesting that meta-learning approaches designed for compositionality can significantly enhance this ability. However, these insights have largely been confined to linguistic problems, leaving their applicability to other tasks an open question. In this study, we extend meta-learning for compositionality to the domain of abstract spatial reasoning. To this end, we introduce Compositional-ARC—a dataset designed to evaluate the capacity of models to systematically generalize from known geometric transformations (e.g., translation, rotation) of abstract two-dimensional objects to novel combinations of these transformations (e.g., translation+rotation). Our results show that a small transformer-based encoder-decoder model, trained via meta-learning for compositionality, can systematically generalize to previously unseen transformation compositions. Notably, despite having only 5.7M parameters, this model significantly outperforms state-of-the-art LLMs—including o3-mini, GPT-4o, and Gemini 2.0 Flash, which fail to exhibit similar systematic behavior—and performs on par with the winning model of the ARC prize 2024, an 8B-parameter LLM trained via test-time training. Our findings highlight the effectiveness of meta-learning in promoting systematicity beyond linguistic tasks, suggesting a promising direction toward more robust and generalizable models.
@@ 1 Introduction
Despite its central role in human cognition, systematic generalization remains a significant challenge in artificial intelligence (Lake & Baroni, 2018; Loula et al., 2018; Hupkes et al., 2020). While large language models have recently demonstrated notable progress across various domains (OpenAI, 2024; Guo et al., 2025), they often fail to combine acquired knowledge in novel scenarios, demonstrating notable difficulties with systematic generalization (Dziri et al., 2023; Ismayilzada et al., 2025; Gendron et al., 2024). The question of whether neural networks can achieve systematicity has been the subject of extensive debate (Fodor & Pylyshyn, 1988; Brakel & Frank, 2009; Calvo & Symons, 2014, inter alia). Recent research by Lake & Baroni (2023) demonstrates that a transformer-based encoder-decoder model, trained via meta-learning for compositionality (MLC), can achieve human-like systematic generalization in processing instructions expressed in a pseudolanguage. By training the model to combine basic units of pseudolanguage into novel sequences over a stream of dynamically changing grammars, Lake & Baroni (2023) show that this model can effectively generalize to previously unseen compositions of language (see Section 2 for further details). While this approach presents a promising direction for addressing systematicity in neural networks, its applicability beyond linguistic contexts remains an open question.
In this study, we extend the MLC framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. Inspired by the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), we introduce Compositional-ARC—a new dataset for assessing systematic generalization in abstract spatial reasoning. Compositional-ARC presents examples of basic geometric transformations (e.g., translation, rotation) applied to abstract two-dimensional objects and tests generalization to previously unseen compositions (e.g., translation+rotation; see Figure 1). Using MLC, we train a small encoder-decoder model on samples from Compositional-ARC and demonstrate that it can systematically generalize to unseen transformation compositions. To the best of our knowledge, this is the first application of MLC to abstract spatial reasoning. In summary, our contributions are:
@@
We introduce Compositional-ARC—a novel dataset, inspired by ARC (Chollet, 2019), that evaluates systematic generalization in abstract spatial reasoning. The dataset includes examples of basic geometric transformations applied to abstract two-dimensional objects and tests generalization to unseen transformation compositions (see Figure 1).
@@
We demonstrate that MLC enables transformer-based models to generalize to unseen compositions of geometric transformations, demonstrating its potential beyond linguistic tasks.
@@
We show that a 5.7M-parameter encoder-decoder model trained via MLC significantly outperforms state-of-the-art general-purpose LLMs such as o3-mini (OpenAI, 2025), GPT-4o (Achiam et al., 2023), and Gemini 2.0 Flash (DeepMind, 2024), which fail to exhibit comparable systematic behavior on Compositional-ARC.
4.
We find that the same MLC model performs on par with the winning model of the ARC Prize 2024, an 8B-parameter LLM trained via test-time training (Franzen et al., 2024).
@@ 2 Background: meta-learning for compositionality
While Lake & Baroni (2023) also evaluate MLC on COGS (Kim & Linzen, 2020) and SCAN (Lake & Baroni, 2018), which test systematic lexical generalization to novel word combinations, their experiments are confined to the linguistic domain. In the following section, we propose Compositional-ARC to show how MLC can be extended to support systematic generalization in abstract spatial reasoning, demonstrating its potential beyond linguistic tasks.
@@ 3 Method
3.1 Compositional-ARC
To test systematicity in abstract spatial reasoning, we leverage the closure property of combined geometric transformations, where the composition of two valid transformations—such as translation, rotation, and reflection—yields another valid geometric transformation (Brannan et al., 2011). Drawing inspiration from the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), we design a task in which abstract objects, defined in a two-dimensional grid environment, are subjected to basic geometric transformations and their compositions (see Figure 1 for examples). We use fixed-size grids, each of which can be represented as a two-dimensional array of integers, where different values correspond to distinct colors. We use integers from 0 to 9, with 0 denoting a black background and the remaining integers mapping to unique colors (see Appendix A.1 for more details). Objects are defined based on color connectivity; that is, each object comprises a group of connected cells sharing the same color. Connectivity is determined by the Moore neighborhood (Bays, 2010), meaning that cells are considered connected if they are directly or diagonally adjacent. Each grid contains either one or two objects. A transformation is represented as a pair of grids, with the input grid displaying the objects before, and the output grid showing them after the geometric transformation. Each transformation affects only one of the objects in the grid. For example, in Figure 1(a), a single L-shaped yellow object is translated one step downward. In Figure 1c, a square blue object in the bottom-right expands toward the neighboring top row. Objects never occlude one another nor extend beyond the boundaries of the grids.
@@ 3.1 Compositional-ARC
To test systematicity, we present few-shot examples of primitive transformations and their level-1 compositions, and evaluate models on previously unseen level-2 compositions. For instance, in Figure 3, models are asked to infer the correct transformation for a previously unseen level-2 composition of indicators, given a set of 12 study examples illustrating primitive transformations and their level-1 compositions. Conceptually, our setup is similar to the few-shot compositional task introduced by Lake & Baroni (2023) (see Section 2), but it replaces the lexical interpretation grammar with a visual interpretation grammar. Specifically, models need to infer which indicator maps to which transformation, and how to compose them to deduce the correct final transformation. For a detailed description of how we algorithmically generate dataset samples, please refer to Appendix A.3.
@@ 3.2 Meta-learning for compositionality in abstract spatial reasoning
To systematically generalize from known geometric transformations to previously unseen transformation compositions, we extend the meta-learning for compositionality (Lake & Baroni, 2023) framework described in Section 2. As in the original MLC approach, we train a transformer-based encoder-decoder model on a dataset of dynamically changing interpretation grammars. However, instead of mapping pseudolinguistic instructions to sequences of abstract symbols, we consider a visual interpretation grammar that associates visual indicators (object shape, color, or proximity to an indicator object) with specific geometric transformations, as described in Section 3.1. An episode in Compositional-ARC is defined as a set of study examples that illustrate the underlying grammar, along with query inputs for which the correct outputs must be inferred. For instance, the episode in Figure 3 contains 12 study examples: six primitive transformations (two per indicator type) and six level-1 compositions (two per composition type). Given the study examples, the model is asked to predict output grids for previously unseen level-2 compositions. By training over a series of episodes with changing visual interpretation grammars, the model needs to abstract and recombine information from the examples in order to predict the correct query transformation composition, as it cannot rely on fixed mappings from indicators to transformations.
@@ 4.1 Task setup
We consider two task setups in this work. The first, denoted as “3-Shot,” is a standard few-shot learning task where models must generate an output grid for a query input that performs a level-2 transformation composition. This prediction is based on three examples illustrating the same level-2 transformation. A visual representation of this setup is provided in Figure 4 in the Appendix. This task evaluates the model’s ability to infer geometric transformations from a limited set of examples.
@@
The second setup, denoted as “Systematicity,” focuses on compositional generalization and differs from the first in the type of few-shot examples presented. As mentioned in Section 3.1, the idea is to test whether models can infer novel compositions from known geometric transformations. To this end, we replace the level-2 few-shot examples with a set of primitive transformations plus level-1 transformation compositions, and query the model to predict the previously unseen level-2 transformation composition, as illustrated in Figure 3. Specifically, we present six primitive transformations—two examples for each indicator (shape-based, color-based, neighbor-based)—and six level-1 transformation compositions, two examples for each level-1 indicator composition (shape+color, shape+neighbor, color+neighbor).
@@
We generate 100,000 episodes, each comprising three few-shot examples for the “3-Shot” task, 12 systematic study examples for the “Systematicity” setup, and ten query input-output grid pairs demonstrating the final level-2 transformation composition. Each episode is characterized by a unique visual interpretation grammar. For instance, in one episode, yellow objects are translated downward by a single cell, while in another, yellow objects are reflected horizontally. To train our encoder-decoder model via MLC, we split the data into 82,908 training, 8,546 validation and 8,546 test episodes. Importantly, the data splits are constructed such that the geometric transformations involved in the final query level-2 compositions differ between the training and evaluation sets. For instance, while the model is trained on basic transformations and a series of transformation compositions (e.g., translation+rotation+reflection), it is tested out-of-distribution on compositions not seen during training (e.g., translation+rotation+extension). For comprehensive statistics of the dataset splits, please refer to Table 5 in the Appendix.
4.2 Large language models
General-purpose LLMs.
@@ 4.2 Large language models
In addition to the model trained via MLC, we evaluate three state-of-the-art general-purpose LLMs on the test set of our proposed dataset: o3-mini (low) (OpenAI, 2025), GPT-4o (Achiam et al., 2023), and Gemini 2.0 Flash (DeepMind, 2024). To textually prompt the models for a given episode, we represent grids as two-dimensional arrays, consistent with prior work (Moskvichev et al., 2023). We also test a multimodal setup in which both an image of the study examples and the input query are provided alongside the text prompt. Due to financial constraints, each model is evaluated on a single test query for each of the 8,546 episodes in the test set. All textual and visual prompts, specific model versions, and decoding parameters are detailed in Appendix C.2.
Domain-specific LLMs.
We further consider two LLMs specifically tailored to ARC-style data: (i) Llama-3.2-3B-ReARC, fine-tuned on the re-ARC dataset (Hodel, 2024)—an extension of 1,000 additional generated examples per sample in ARC—and (ii) Mistral-NeMO-Minitron-8B-Full, trained on a broad range of ARC-style data, including re-ARC, Concept-ARC (Moskvichev et al., 2023), and ARC-Heavy (Li et al., 2025). These models were proposed by Franzen et al. (2024) and placed 1st in the ARC prize 2024.111https://arcprize.org/competitions/2024 Note that in addition to fine-tuning, these models use an ARC-customized tokenizer, extensive data augmentation during training and inference, a generation procedure that leverages depth-first search to produce multiple solution candidates, and a refined candidate-selection step. The authors also employ test-time training (TTT), which further fine-tunes models on the few-shot input–output grid pairs from the final test set. We use both models with their default parameters. For additional details, please refer to the original paper (Franzen et al., 2024) or Appendix C.2.
@@ 4.3 Evaluation metrics
To evaluate the quality of the generated output grids, we use three different metrics: i) exact match accuracy, ii) color accuracy, and iii) shape accuracy. Exact match accuracy requires that a prediction is accurate only if every cell matches the target grid. Color accuracy checks whether predicted objects match target colors, ignoring shape and location. Shape accuracy checks whether predicted objects match target shapes, ignoring color and location. Formal definitions are provided in Appendix C.1.
| Model | Exact Match Accuracy [%] | Color Accuracy [%] | Shape Accuracy [%] | |
3-Shot |
GPT-4o | 22.28 | 99.67 | 57.02 |
| + image | 19.42 | 99.75 | 54.56 | |
| Gemini 2.0 Flash | 30.08 | 99.92 | 52.34 | |
| + image | 17.19 | 99.79 | 35.86 | |
| o3-mini (low) | 64.04 | 99.89 | 68.74 | |
| Llama-3.2-3B-ReARC | 85.85 | 98.57 | 86.05 | |
| Mistral-NeMO-Minitron-8B-Full | 95.71 | 99.85 | 96.78 | |
| MLC (ours) | 99.92 | 100.00 | 99.92 | |
Systematicity |
GPT-4o | 0.99 | 99.23 | 9.82 |
| + image | 0.86 | 97.94 | 7.50 | |
| Gemini 2.0 Flash | 2.66 | 99.68 | 12.81 | |
| + image | 2.05 | 99.28 | 9.60 | |
| o3-mini (low) | 0.53 | 99.10 | 5.65 | |
| Llama-3.2-3B-ReARC | 0.87 | 99.94 | 2.54 | |
| + test-time training | 73.70 | 100.00 | 86.88 | |
| Mistral-NeMO-Minitron-8B-Full | 0.70 | 99.99 | 9.75 | |
| + test-time training | 78.20 | 100.00 | 88.26 | |
| MLC (ours) | 78.26 | 97.88 | 80.49 |
@@ 5 Results
In Table 1, we report the performance of the model trained via MLC, alongside the LLMs we evaluate on the two task setups, as described in Section 4.1.
@@
We begin by examining model performance on the “3-Shot” task, where models are given three input-output examples illustrating the final transformation composition (see Figure 4 in the Appendix). Despite this guidance and the relatively simple transformations involved, general-purpose LLMs such as GPT-4o and Gemini 2.0 Flash struggle with the task: GPT-4o reaches an accuracy of only 22.28%, while Gemini 2.0 Flash performs slightly better at 30.08%. The long-chain-of-thought model o3-mini achieves a modest accuracy of 64.04%. In contrast, domain-specific models such as Llama-3.2-3B-ReARC, and Mistral-NeMO-Minitron-8B-Full perform significantly better. While Llama-3.2-3B-ReARC achieves an accuracy of 85.85%, Mistral-NeMO-Minitron-8B-Full reaches up to 95.71%. Note that we do not employ test-time training in this setup, as it would contradict the out-of-distribution test setup described in Section 4.1. Notably, the 5.7M-parameter encoder-decoder model trained via MLC outperforms both general-purpose and domain-specific LLMs, with an accuracy of 99.92%, despite having only a fraction of the parameters. We further find that all models predict object color nearly perfectly. For GPT-4o and Gemini 2.0 Flash, we observe that shape accuracy is significantly higher than exact match accuracy. This discrepancy suggests that while these models are often able to predict the correct shape and color of an object, they frequently fail to accurately predict its final position. Interestingly, both models show lower accuracy when visual input is added to the textual prompt, likely due to modality alignment challenges (Masry et al., 2025) or limitations in leveraging the visual content for reasoning.
@@
In the “Systematicity” task, models are asked to infer the correct final transformation composition from a set of study examples that represent more basic, decomposed transformations (see Figure 3 for an example). As shown in Table 1, all general-purpose LLMs perform poorly on this task. For instance, GPT-4o achieves an accuracy of 0.99%, while Gemini 2.0 Flash reaches 2.66%. Interestingly, o3-mini, the best-performing general-purpose model on the “3-Shot” task, performs worst in this setting, with an accuracy of only 0.53%. For the domain-specific LLMs, we find that test-time training (TTT)—where models are additionally fine-tuned on the study examples’ input-output grid pairs of the test set—significantly improves performance. While Llama-3.2-3B-ReARC achieves only 0.70% accuracy without TTT, performance increases to 73.70% with TTT. Similarly, Mistral-NeMO-Minitron-8B-Full’s accuracy increases from 0.70% to 78.20% with TTT. We hypothesize that training on the systematic study examples of the test data (demonstrating primitive and level-1 transformations) teaches the models how to abstract and compose transformations for the final input query. We further find that the much smaller 5.7M-parameter MLC model performs on par with the domain-specific LLMs trained via TTT, slightly outperforming Mistral-NeMO-Minitron-8B-Full with an accuracy of 78.26%. Importantly, as described in Section 4.1, the MLC model has never seen the specific level-2 compositions of the test data during training, but was instead optimized on a distinct set of transformation compositions (see data split for seed 1860; Table 5 in the Appendix). Consistent with our findings from the 3-shot learning task, models generally succeed in predicting the correct object colors. However, shape accuracy declines markedly. A qualitative example of the models’ predictions is shown in Figure 3, with additional examples in Figures 7– 8 in the Appendix. The strong performance of the small MLC model highlights the effectiveness of this training strategy in promoting systematic generalization to novel transformation compositions. The model not only learns to infer a visual interpretation grammar from a limited number of study examples but also generalizes to novel transformation compositions that it has never encountered during training.
| Model | Exact Match Accuracy [%] | Color Accuracy [%] | Shape Accuracy [%] |
| MLC (3-Shot) | 98.78 1.99 | 100.00 0.00 | 98.79 1.98 |
| MLC (Systematicity) | 86.73 6.03 | 99.36 0.70 | 87.55 5.45 |
| - no copy task | 69.05 9.23 | 99.43 0.38 | 70.60 9.23 |
| - no primitives | 75.27 12.95 | 99.56 0.50 | 76.92 11.23 |
| - no level-1 compositions | 21.01 19.07 | 94.72 7.41 | 23.03 19.08 |